### v4

# Deflickering Vision-Based Occupancy Networks through Lightweight Spatio-Temporal Correlation

Abstract
Vision-based occupancy networks (VONs) provide an end-to-end solution for reconstructing 3D environments in autonomous driving. However, existing methods often suffer from temporal inconsistencies, manifesting as flickering effects that compromise visual experience and adversely affect decision-making. While recent approaches have incorporated historical data to mitigate the issue, they often incur high computational costs and may introduce noisy information that interferes with object detection. We propose OccLinker, a novel plugin framework designed to seamlessly integrate with existing VONs for boosting performance. Our method efficiently consolidates historical static and motion cues, learns sparse latent correlations with current features through a dual cross-attention mechanism, and produces correction occupancy components to refine the base network’s predictions. We propose a new temporal consistency metric to quantitatively identify flickering effects. Extensive experiments on two benchmark datasets demonstrate that our method delivers superior performance with negligible computational overhead, while effectively eliminating flickering artifacts.
I Introduction
Vision-based Occupancy Networks (VONs) [OccFormer, zhang2025occnerf, VoxFormer, mei2024camera, yang2024adaptiveocc, jiang2024symphonize] have emerged as a powerful paradigm for reconstructing surrounding environments from ego-centric multi-view images. Typically, VONs represent the scene as a structured grid of voxels [Occ3D] by learning accurate mappings from 2D visual cues to 3D occupancy representations. This approach enables the extraction of both geometric structure and semantic understanding, thereby facilitating various applications, including indoor navigation [liu2024volumetric, li2023stereovoxelnet] and autonomous driving [xu2025exploiting, xu2025temporal, ouyang2024linkocc, yang2024adaptiveocc].
Despite these progresses, the occupancy results generated by existing VONs often exhibit inaccuracies and visual imperfections. A particularly problematic phenomenon, which we term flickering, manifests as instability in the constructed scenes: objects may appear or disappear abruptly, accompanied by various artifacts. This lack of temporal consistency and accuracy not only confuses the driving control system but also seriously degrades the visual experience for human drivers. The primary causes of flickering include sensor noise, occlusion, model limitations, and thresholding issues. As illustrated in Fig. 1, a pedestrian occluded by a tree provides incomplete visual cues, posing a significant challenge to the prediction models. In this case, SurroundOcc [surroundOcc], which operates in a frame-by-frame “detection” paradigm, fails to detect the pedestrian in Frame #24. The temporary absence of the person will create a flickering effect when the results are displayed to the user.
Fig. 1: Enhancing existing VONs with OccLinker. (i) In the upper part, the first row shows three consecutive images where the ego-vehicle moves forward along the right side of the road. The blue box highlights a pedestrian as a sparse visual cue, and the second row provides the corresponding zoomed-in views. (ii) In the lower part, we compare different types of VON methods. The first row shows a standard 3D VON [surroundOcc] that ignores historical information, resulting in the pedestrian being missed in certain frames. The second row presents a history-aware VON [bevdet4d] that utilizes past frames but still produces incomplete voxels due to suboptimal temporal association. In contrast, our module, when integrated with the 3D VON [surroundOcc], achieves smoother and more complete occupancy predictions through effective spatio-temporal correlation.
An effective approach to mitigate the temporal inconsistency problem is to incorporate historical information when predicting the current occupancy. Recent history-aware VONs [opus, bevdet4d, SparseOcc_Liu, viewformer] have explored this direction and can be broadly categorized into two classes: dense Bird’s Eye View (BEV)/voxel-based temporal fusion methods and sparse query or representation based methods. Specifically:
  • Dense methods [bevdet4d, fb_occ, panoocc, viewformer, chen2025rethinking, DriveWorld, ye2024cvt] perform dense temporal modeling, typically by propagating, aligning, or aggregating historical BEV features or volumetric representations into the current frame. These approaches incur substantial computational and memory costs, often limiting the feasible temporal window or spatial resolution. For example, as shown in Fig. 2, BEVDet4D-Occ produces the largest inference latency. Moreover, dense temporal aggregation may propagate redundant or noisy features across broad spatial regions, leading to unstable predictions in dynamic areas. As illustrated in the lower part of Fig. 1, the pedestrian appears unclear in the BEVDet4D-Occ results, likely due to coarse-grained temporal fusion.
  • Sparse methods seek to reduce the complexity of dense voxel fusion by operating on sparse queries [opus, oh20253d] or representations [SparseOcc_Liu, liao2025stcocc, jiang2025gausstr]. By focusing computation on most relevant latent regions, methods such as SparseOcc and the OPUS series typically achieve lower inference latency (see Fig. 2). These methods often require multi-stage recovery or decoding processes to reconstruct dense occupancy. For instance, SparseOcc employs diffusion-based propagation combined with feature pyramids to densify non-empty regions across scales, whereas OPUS adopts a coarse-to-fine cross-attention mechanism to progressively decode sparse queries into 3D space. These designs increase memory usage and produce of reduced voxel-level accuracy.
Fig. 2: Comparison with history-aware VON methods in terms of mIoU and running costs. Points closer to the bottom-left indicate better efficiency, with lower memory usage and faster inference. Larger circles denote higher prediction quality with mIoU shown in numbers. Our method, when integrated with ViewFormer, achieves the best mIoU and the lowest memory consumption, while maintaining competitive inference latency.
Different from existing methods, we explore an alternative approach: enhancing temporal consistency through a lightweight plug-in module that leverages historical information without redesigning the backbone, altering the fusion strategy, or retraining the base VON from scratch. Specifically, our approach only requires training a lightweight module on top of existing VON predictions or intermediate features, making it both practical and broadly applicable.
To this end, we propose a novel method, termed OccLinker. Our approach improves occupancy prediction by discriminating and extracting relevant motion and static cues from historical frames, emphasizing lightweight yet effective spatio-temporal correlation modeling. OccLinker follows a three-stage framework. First, it converts historical static and motion cues into queries through sparse tokenization. Next, an independent latent correlation module is introduced to attentively associate current features with historical cues within a shared latent space. Finally, the system generates correction occupancy that enables refinement of the base network’s results. Our method offers two key advantages:
  • Seamless integration with state-of-the-art (SOTA) VONs for significantly enhanced prediction accuracy. When integrated with SurroundOcc [surroundOcc] on its benchmark, it elevates IoU and temporal consistency of moving objects (as defined in the Experiment section) by 1.63% and 2.31, respectively. Similarly, when combined with ViewFormer [viewformer] on the Occ3D [Occ3D] benchmark, it achieves gains of 0.24% in IoU and 2.87 in temporal consistency;
  • Low cost through lightweight correlation operation, which surpasses recent history-aware VONs [opus, bevdet4d, fb_occ] with much lower computational costs. For instance, OPUS-L [opus] requires 10.33 GB of GPU memory and 0.16 seconds per frame for inference. In contrast, ViewFormer+OccLinker only needs 4.58 GB memory and 0.12 seconds. Notably, the OccLinker plug-in itself uses just 105 MB of memory and adds only 0.01 seconds of latency.
These two advantages jointly push our method beyond the usual accuracy‑efficiency Pareto frontier. As illustrated in Fig. 2, existing history‑aware approaches demonstrate a clear trade-off between performance and cost. In contrast, our method, when combined with ViewFormer, achieves superior accuracy without sacrificing efficiency.
In summary, the contributions of this paper are three-fold:
  • We introduce OccLinker, a parameter-efficient plug-in that integrates seamlessly with existing 3D VONs without requiring any modifications to the base network. It incurs only minimal GPU memory overhead and negligible inference latency, while consistently enhancing prediction accuracy.
  • We propose a lightweight spatio-temporal correlation learning method that effectively fuses historical cues, sparse motion features, as well as current static tokens. These cues are mapped into a shared compact latent space, enabling pixel-to-voxel correlation through adaptive 3D occupancy aggregation.
  • We devise a new temporal consistency metric to help analyze the effectiveness of various occupancy methods. Extensive experiments are conducted on both SurroundOcc [surroundOcc] and Occ3D [Occ3D] benchmarks, showing the of advantages of our approach in accuracy while maintaining real-time inference efficiency.
The remainder of this article is structured as follows: Section II reviews related work; Section III elaborates on the design of our proposed method; Section IV presents experimental results; and Section V concludes this paper with a discussion of future work.
II Related Work
II-A 3D Scene Reconstruction
3D scene reconstruction technology has evolved rapidly in autonomous driving. Early studies focused on mathematical methods [MonoScene] and camera parameters [3D_reconstruction_camera_para_1, 3D_reconstruction_camera_para_2, 3D_reconstruction_camera_para_3] to facilitate 3D scene reconstruction from 2D inputs. MonoScene [MonoScene] explored the path of scene reconstruction from 2D to 3D using monocular RGB images, setting a groundbreaking benchmark. Later, more learning methods were applied. 1) Multi-view reconstruction [Multi_View_reconstruction_1, Multi_View_reconstruction_2, shi2024occupancy], which integrates multiple perspectives for enhanced depth estimation. 2) Multimodal methods combining RGB with depth [multimodal_rgb_depth_1, multimodal_rgb_depth_2] or RGB with LiDAR [multimodal_rgb_lidar_1] data, improving the robustness and precision of reconstructions.
Fig. 3: Overview of OccLinker. A frozen base 3D VON [surroundOcc, MonoScene, viewformer] runs at a fixed frequency to extract static texture features by $\Phi_{\rm SE}^{*}$ and generate initial occupancy prediction by $\Phi_{\rm SD}^{*}$. We employ $\Phi_{\rm ME}$ for extracting motion features of intermediate frame differences, and OccLinker uses the static and motion features from the recent temporal window as queries, and the current static feature as the key and value. Through lightweight tokenization encoder $\phi$ and dual cross-attention module $\Theta$, it constructs spatio-temporal correlations and outputs a correction term $\Delta O_{t}$ to refine the initial prediction.
II-B 3D VONs
3D VONs extend the capabilities of 3D scene reconstruction technologies to enhance dynamic and precise environmental perception for autonomous vehicles. This task involves learning the associations between pixel-level visual cues and the occupancy states of 3D voxels. Recent approaches often employ Transformers [attention_is_all_you_need] to effectively model these complex relationships, achieving significant advancements in accuracy and efficiency. 3D occupancy prediction have utilized multiple data inputs such as LiDAR [Panoptic_PolarNet, PointOcc, li2025occmamba] and depth data [S3CNet_depth, Li_2019_CVPR_depth, Li_2020_CVPR_depth, See_and_think_depth, Li_Zou_Li_Zhao_Gao_2020_depth, Depth_Based_Semantic_Scene_depth, Two_Stream_3D_Semantic_Scene_Completion_depth] to predict voxel occupancy. However, with advancements in camera technology and image processing, current research predominantly adopts an image-based end-to-end approach [surroundOcc, TPVFormer, VoxFormer, selfocc, RenderOcc]. These modern methods use data from single frames to predict occupancy, focusing on integrating rich semantic information at lower costs. However, due to the high-dimensional visual input and dense grid-based output, end-to-end methods suffer from significant computational intensity. To address this, more recent works [jiang2025gausstr, huang2024gaussianformer, huang2025gaussianformer, gan2025gaussianocc, zhu2025voxelsplat] have incorporated Gaussian point clouds as an intermediate representation. These point clouds are projected into the 3D occupancy space using splatting techniques, which significantly reduce memory consumption.
II-C History-aware VONs
Recent advances in spatio-temporal perception have extended 3D VONs into temporal domains, primarily through two paradigms: (i) 4D VONs that estimate current and future states (e.g., Cam4DOcc [Cam4docc], OccSora [OccSora]), and (ii) history-aware VONs [fb_occ, SparseOcc_Liu, bevdet4d, opus] leverage historical features to refine current-frame predictions. In this paper, we focus primarily on the second category, which, as discussed in the Introduction, can be further classified into dense BEV or voxel-based temporal fusion [bevdet4d, fb_occ, panoocc, viewformer, chen2025rethinking, DriveWorld, ye2024cvt] and sparse query [opus, oh20253d] or representation-based methods [SparseOcc_Liu, liao2025stcocc, jiang2025gausstr]. For instance, the sparse method OPUS [opus] uses a learnable query-based encoder-decoder architecture and retains coarse-grained concatenation of raw image features during early fusion stages, while the dense method BEVDet4D-Occ [bevdet4d] performs spatio-temporal alignment from a BEV perspective. In contrast, we propose a different approach, using a lightweight plugin that effectively discriminates and extracts relevant motion-static attention, complementing the current occupancy predictions. Our method does not conflict with existing history-aware approaches and can seamlessly integrate with them or other 3D VONs to enhance both occupancy accuracy and temporal consistency.
III Methodology
III-A Problem Formulation
In the context of vision-based occupancy networks (VONs) [xu2025exploiting, VoxFormer, surroundOcc, yang2024adaptiveocc, ouyang2024linkocc] for autonomous driving, the ego-vehicle typically collects two types of data: (i) multi-view RGB frames from surround-view cameras (typically six), captured at a relatively low and fixed frequency – we refer to these as keyframes; (ii) higher-frequency motion information between keyframes, such as optical flow or frame differences.
As illustrated in Fig. 3, given a temporal window $\mathcal{W}$ of length $L$, the input includes a sequence of $L$ historical keyframes denoted by $S_{t-L:t-1}$, one current keyframe $S_{t}$, and a corresponding sequence of $L$ motion cues $M_{t-L:t-1}$, where each motion cue represents frame differences over a short interval $\Delta l$. Let $|V|$ denote the number of cameras. For simplicity, we merge all multi-view inputs into a unified notation and do not distinguish between individual camera views. Both RGB and motion frames are assumed to have the same spatial resolution.
Typically, a base 3D VON $\Phi$ first encodes the current keyframe $S_{t}$ using a 2D encoder $\Phi_{{\rm SE}}$ to extract static appearance features $s_{t}$. These features are then lifted into 3D space via a 2D-to-3D decoder $\Phi_{{\rm SD}}$, resulting in an initial occupancy prediction $O_{t}^{\prime}$. That is:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{t}=\Phi_{{\rm SE}}^{*}(S_{t}),\quad O_{t}^{\prime}=\Phi_{{\rm SD}}^{*}(s_{t}),$} (1)
where $\cdot^{*}$ indicates frozen weights from the pre-trained foundational network. Since the base VON operates in a frame-by-frame manner, the historical static features $s_{t-L:t}$ extracted by the frozen static encoder $\Phi_{\rm SE}^{*}$ are pre-stored and reused for temporal correlation modeling.
Our goal is to design a lightweight plug-in module, parameterized by $\Theta$, that associates the motion features within a temporal window of length $L$ with the static features $s_{t-L:t}$ extracted by the base VON, thereby enhancing the foundational prediction $O_{t}^{\prime}$ at each timestep $t$. To achieve this, the intermediate frame differences, denoted by $M_{t-L:t-1}$, is first encoded into a latent representation using a motion encoder $\Phi_{\rm ME}$, yielding temporal motion features:
$m_{t-L:t-1}=\Phi_{\rm ME}(M_{t-L:t-1}).$ (2)
To enable efficient feature linking, we employ a lightweight tokenization encoder $\phi$ that transforms both the static features $s_{t-L:t}$ and the motion features $m_{t-L:t-1}$ into a compact sparse token set. Subsequently, the cross-attention module $\Theta$ establishes spatio-temporal correlations within this latent token space, enhancing prediction smoothness and accuracy while compensating for missing voxels in the current occupancy estimate. The resulting enhancement component $\Delta O_{t}$ is formulated as:
$\Delta O_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Theta}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}(s_{t-L:t}),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}(m_{t-L:t-1})).$ (3)
Then, the final enhanced occupancy prediction becomes:
$\hat{O_{t}}={\rm Softmax}(O_{t}^{\prime}+\Delta O_{t}).$ (4)
Here, $O_{t}^{\prime}$ and $\Delta O_{t}$ denote the occupancy logits, each with a shape of $C\times H\times W\times D$, where $C$ is the number of semantic occupancy classes, and $H$, $W$, and $D$ correspond to the height, width, and depth (length) of the 3D occupancy data, respectively. The ${\rm Softmax}(\cdot)$ function is applied along the class dimension to convert the logits into class probabilities.
Notably, our method focuses on enhancing the current-frame occupancy using historical and motion cues, rather than forecasting future occupancy. The design principle emphasizes lightweight integration without retraining or modifying the base network. Fig. 3 presents the overall design of our OccLinker. In contrast to previous approaches [ouyang2024linkocc, viewformer, fb_occ, SparseOcc_Liu, bevdet4d, opus] that depend on redundant historical data, OccLinker compute the enhancement component $\Delta O_{t}$ using a dual cross-attention mechanism [attention_is_all_you_need], designed to distill essential cues from both static and motion features while capturing compact latent correlations. First, we apply a sparse tokenization strategy to transform $s_{t-L:t}$ and $m_{t-L:t}$ into two separate sets of queries, as described in Section III-B. Then, we establish spatio-temporal correlation in a compact shared latent space, detailed in Section III-C. Finally, the optimization procedure is described in Section III-D, and the pseudo-code for the OccLinker inference pipeline is summarized in Algorithm 1.
Input: Keyframes $S_{t-L:t}$ and motion cues $M_{t-L:t}$; Frozen static encoder $\Phi_{\rm SE}^{*}$ and decoder $\Phi_{\rm SD}^{*}$ from the base VON; Trainable motion encoder $\Phi_{\rm ME}$ and trainable components in OccLinker, including tokenization encoder $\phi$ and dual cross-attention module $\Theta$.
Output: Enhanced occupancy prediction $\hat{O}_{t}$.
1 Initialize temporal context window $\mathcal{W}\leftarrow\emptyset$, $|\mathcal{W}|=L$;
2 while True do
 // Stage 1: Base VON Inference
3   Extract static features from the current frame: $s_{t}\leftarrow\Phi_{\rm SE}^{*}(S_{t})$;
4   Predict initial occupancy logits: $O_{t}^{\prime}\leftarrow\Phi_{\rm SD}^{*}(s_{t})$;
 // Stage 2: Spatio-Temporal Linking
5   Extract motion features from the high-frequency motion cues: $m_{t}\leftarrow\Phi_{\rm ME}(M_{t})$;
6   Append $(s_{t},m_{t})$ to the temporal window $\mathcal{W}$ and discard the oldest entry if $|\mathcal{W}|>L$;
7   Retrieve recent static and motion features $s_{t-L:t-1}$, $m_{t-L:t-1}$, and $s_{t}$ from $\mathcal{W}$;
8   Convert all retrieved features into sparse tokens via Eq. (9);
9   Establish spatio-temporal correlations within these tokens via Eq. (10);
10   Decode the attention outputs into a correction component via Eq. (11);
 // Stage 3: Occupancy Refinement
11   Refine the initial occupancy prediction via Eq. (4);
12 Advance to the next timestep: $t\leftarrow t+1$;
13
14 end while
Algorithm 1 OccLinker inference pipeline
III-B Sparse tokenization for query construction
To effectively extract static and motion cues for downstream correlation modeling, we first encode the raw inputs into compact token representations. This process is composed of two parallel modules: a static feature encoder (SE) and a motion feature encoder (ME), followed by a lightweight tokenization encoder (TE).
Static feature encoder
We extract static spatial features from both the current and previous keyframes by directly reusing the 2D vision backbone modules, denoted as $\Phi_{\rm SE}$, from foundational vision-based occupancy networks (e.g., SurroundOcc [surroundOcc] or ViewFormer [viewformer]). Let $S_{t}$ represent the current surround-view image and $S_{t-L:t-1}$ denotes the historical keyframes within a temporal window of size $L$. The corresponding static features are computed as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{t-L:t}=\Phi_{\rm SE}^{*}(S_{t-L:t})$}, (5)
where $\Phi_{\rm SE}^{*}$ shares parameters across all time steps.
Motion feature encoder
Although occupancy prediction is typically performed using keyframes sampled at a fixed rate, the ego-vehicle continuously captures high-frequency RGB images. These intermediate frames, located between consecutive keyframes, contain rich motion cues that are valuable for spatial-temporal perception. To efficiently extract motion information, we adopt a simple yet effective frame-difference technique [Frame_difference]. For each pair of adjacent keyframes $(S_{i},S_{i+1})$ within the historical window [tL,t][t-L,t], where $t-L\leq i<i+1\leq t$. We compute the difference between any two intermediate frames indexed by $(\tau_{a},\tau_{b})$ satisfying $i\leq\tau_{a}<\tau_{b}\leq i+1$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{i:i+1}=\{S^{\tau_{b}}_{i:i+1}-S^{\tau_{a}}_{i:i+1}\}$}. (6)
All such differences across the $L$ intervals are concatenated as the motion input:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M_{t-L:t-1}=\text{Concat}\left(\left\{F_{i:i+1}\mid i\in[t-L,t-1]\right\}\right)$}, (7)
where $\text{Concat}(\cdot)$ denotes the concatenation operation performed along the temporal dimension, combining motion cues from all intervals in the historical window.
The resulting difference images $M_{t-L:t}$ are then used as input to the motion encoding module. Since these frame difference images typically share the same resolution as the original RGB images – often high-resolution [nuScenes] (e.g., $1600\times 900$) – direct processing can be expensive and unsuitable for real-time applications. To address this, we first downsample $M_{t-L:t-1}$ by a factor of five to reduce the spatial dimensions. Moreover, unlike RGB images that contain complex textures and color variations, frame difference images primarily capture motion boundaries. Therefore, we employ a lightweight three-layer convolutional neural network (CNN) followed by batch normalization (BN) to efficiently extract motion cues. This network is denoted by $\Phi_{\rm ME}$. Formally, the resulting motion features are computed as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m_{t-L:t-1}=\Phi_{\rm ME}(M_{t-L:t-1}).$} (8)
Sparse tokenization encoder
The extracted feature $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ are then fed into a shared tokenization encoder that transforms them into compact token sets. Specifically, a $1{\times}1$ convolution is applied to first reduce channel dimensions, followed by a spatial unfolding operation that divides each image into $p{\times}p$ patches. Within each patch, pixel-wise features are averaged to form a $v$-dimensional vector. This process is applied to each camera view independently and results in the following tokenization operation $\phi$:
$\phi(I)=\overline{\mathcal{U}(p,p)(I\circledast\Theta_{1\times 1})},$ (9)
[…]
The final output $E$ is a set of sparse tokens with shape $|V|\cdot\lfloor\frac{h}{p}\rfloor\cdot\lfloor\frac{w}{p}\rfloor,1,d_{token}]$, where each token retains localized spatial or motion semantics. This tokenized format is especially well-suited for attention-based fusion in our multi-stream integration module, as discussed next.
III-C Spatio-temporal correlation strategy
To efficiently exploit historical static and motion cues while avoiding redundant computation, we adopt a spatio-temporal correlation strategy based on dual multi-head attention operations. Unlike previous approaches that directly fuse all temporal inputs indiscriminately [bevdet4d, yang2024adaptiveocc], our method disentangles the correlation modeling process into two independent streams: one for static context and another for motion cues. These correlations are computed in a shared compact latent space, improving both interpretability and efficiency.
Cross-attention-based spatio-temporal reasoning
Given the sparse tokens $s_{t}$, $s_{t-L:t-1}$, and $m_{t-L:t-1}$ extracted in Section III-B, we use $s_{t}$ (the current frame token) as the reference context. The goal is to model how past appearance and motion features relate to the current frame through dual multi-head attention mechanisms (MHAM) [attention_is_all_you_need].
  • Historical static attention: We use the tokenized historical static features $\phi(s_{t-L:t-1})$ as queries ($Q_{\rm sta}$), while the tokenized current feature $\phi(s_{t})$ serves as both keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). Each component is projected into a shared latent space via independent learnable transformations. This MHAM produces the output $f_{\rm sta}$, capturing long-range spatial appearance correlations, as illustrated in the first column of Fig. 4.
  • Historical motion attention: Similarly, the motion tokens $\phi(m_{t-L:t-1})$ are used as queries ($Q_{\rm mot}$), and the same current feature tokens $\phi(s_{t})$ are reused as keys ($K_{\rm cur}$) and values ($V_{\rm cur}$). The resulting attention output $f_{\rm mot}$ captures temporally aligned motion dynamics, as shown in the second column of Fig. 4.
Each attention stream is formulated as:
$f_{\rm sta}$ $=\mathrm{Softmax}\left(\frac{W_{\rm sta}^{Q}Q_{\rm sta}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$ (10)
$f_{\rm mot}$ $=\mathrm{Softmax}\left(\frac{W_{\rm mot}^{Q}Q_{\rm mot}\left(W_{\rm cur}^{K}K_{\rm cur}\right)^{\top}}{\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}}\right)\left(W_{\rm cur}^{V}V_{\rm cur}\right),$
where $W^{Q}_{*}$, $W^{K}_{*}$, and $W^{V}_{*}$ are independent learnable projection matrices for each stream. The $\mathrm{Softmax}(\cdot)$ function is used to normalize the attention scores, and the scaling factor $\sqrt{d_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}grad}}}$ helps stabilize gradients during training [attention_is_all_you_need].
Fig. 4: Correlation pipeline in OccLinker. Using $L=1$ as an example, OccLinker explicitly separates static and dynamic correlations by treating current static features as keys and values, while using historical static features and intermediate motion features as queries.
Feature decoding and correction
The attention outputs $f_{\rm sta}$ and $f_{\rm mot}$, together with the current token $f_{\rm cur}=K_{\rm cur}$, are passed through separate lightweight decoding heads to reconstruct the spatio-temporal correction signal $\Delta O_{t}$. Each stream is first reshaped using 3D deconvolution layers followed by ReLU activations to recover the spatial dimensions. The decoding network is parameterized by $\Theta_{\rm D}$. This produces three correction results: $\Delta O_{\rm sta}$ from static attention, $\Delta O_{\rm mot}$ from motion attention, and $\Delta O_{\rm cur}$ from the current frame as the baseline. These results are then concatenated along the feature channel and aggregated using an additional 3D convolutional layer:
$\Delta O_{t}={\rm Concat}(\{\Delta O_{\rm sta},\Delta O_{\rm mot},\Delta O_{\rm cur}\})\circledast\Theta_{3\times 3\times 3},$ (11)
where $\Theta_{3\times 3\times 3}$ is a learnable 3D convolution kernel.
The fused correction $\Delta O_{t}$ is then added to the initial occupancy logits $O_{t}^{\prime}$ from the base model, as defined in Eq. (4), yielding the final enhanced prediction $\hat{O}_{t}$. This integration augments the base prediction with temporally aligned static and motion-aware cues, while maintaining full compatibility with any existing foundational VONs.
The proposed correlation design improves both interpretability and efficiency by explicitly separating static and dynamic contexts, enabling robust enhancement without modifying the backbone network.
Notably, we do not rely on any explicit coordinate transformation between frame $t$ and historical frames tL:t1t-L:t-1. Unlike many occupancy prediction methods [xu2025exploiting, viewformer] that incorporate ego-motion information (e.g., relative translation and rotation from IMU or odometry) for geometric alignment or implicit motion compensation, OccLinker directly associates historical static and dynamic cues in the sparse latent space. This design keeps OccLinker lightweight, sensor-agnostic, and fully compatible with both single-frame and history-aware VONs, without requiring additional motion supervision or modifying their original coordinate systems.
III-D Optimization
As shown in Fig. 3, the static encoder $\Phi_{\rm SE}$ is directly inherited from well-established VON architectures [MonoScene, surroundOcc, viewformer]. In contrast, the motion encoder $\Phi_{\rm ME}$ is a lightweight, trainable module specifically designed to extract temporal dynamics and compensate for missing or incomplete voxel predictions caused by occlusions or visual ambiguity. In scenarios where the static encoder $\Phi_{\rm SE}$ produces suboptimal outputs, $\Phi_{\rm ME}$ serves as a complementary component by injecting motion-aware corrections. Notably, the motion encoder is unlikely to dominate or collapse unless the frozen static features are already sufficiently complete to support accurate occupancy predictions on their own.
To maintain stability and efficiency during training, we freeze the static encoder $\Phi_{\rm SE}$ and only optimize the proposed enhancement module OccLinker along with the motion encoder $\Phi_{\rm ME}$. This design yields two key advantages: (1) it avoids potential gradient interference in $\Phi_{\rm SE}$, ensuring more stable optimization; and (2) it accelerates convergence of our plug-in module by leveraging reliable static features extracted by the pre-trained backbone.
To ensure a fair comparison and retain consistency with prior work, we adopt the original loss functions used by each benchmark during training. For the SurroundOcc [surroundOcc] benchmark, we employ the following composite loss function:
$\mathcal{L}=\mathcal{L}_{\rm ce}+\mathcal{L}_{\rm sem}+\mathcal{L}_{\rm geo},$ (12)
where $\mathcal{L}_{\rm ce}$ is a voxel-wise cross-entropy loss for occupancy classification [cross_entropy_loss_2], $\mathcal{L}_{\rm sem}$ enforces semantic consistency across space [Semantic_loss], $\mathcal{L}_{\rm geo}$ encourages geometric coherence between neighboring voxels [geometry_affinity_loss].
For the Occ3D [Occ3D] benchmark, following the convention in ViewFormer [viewformer], we apply the following loss:
$\mathcal{L}=\mathcal{L}_{\rm focal}+\mathcal{L}_{\rm ce}+\mathcal{L}_{\rm ls}+\lambda\mathcal{L}_{\rm l1},$ (13)
where $\mathcal{L}_{\rm focal}$ is the focal loss [focal_loss] for handling class imbalance, $\mathcal{L}_{\rm ls}$ is the lovasz softmax loss [lovasz_loss] for improving boundary-level segmentation, $\mathcal{L}_{\rm l1}$ is an L1 regression loss (scaled by $\lambda$) to supervise motion-induced voxel correction.
IV Experiments
IV-A Experimental settings
Benchmark
[…]
Implementation details
Both the SurroundOcc and Occ3D benchmarks adopt the same 3D occupancy resolution of $(H,W,D)=(200,200,16)$ and utilize the same number of surround-view cameras, with $|V|=6$. However, they differ in semantic granularity: SurroundOcc defines $C=17$ semantic classes, while Occ3D provides $C=18$ for finer-grained occupancy labeling. Following the respective conventions [surroundOcc, viewformer] for static feature dimensions, we adopt a feature map resolution of $(h,w)=(116,200)$ for SurroundOcc and $(h,w)=(32,88)$ for Occ3D. Notably, both $s_{t-L:t}$ and $m_{t-L:t}$ share the same dimensionality of $|V|\times c\times w\times h$.
We employ the $\Phi_{\rm ME}$ module with a temporal sampling interval of $\Delta l\approx 0.1$ seconds to extract motion features from raw nuScenes images at a resolution of $900\times 1600$. Within the $\Phi_{\rm ME}$ module, input images are first downsampled by a factor of 5, resulting in $180\times 320$ feature maps to reduce GPU memory consumption. Motion inputs are computed using either frame difference operations [Frame_difference] or optical flow analysis [OpticalFlow]. The token dimension and the gradient scaling factor are both set to $d_{token}=d_{grad}=32$, with the number of temporal patches set as $p=6$. For the convolutional layer $\Theta_{1\times 1}$, the kernel size, stride, padding, input channels, and output channels are set to $(1,1)$, 11, 0, 512, and $d_{token}$, respectively. In contrast, for $\Theta_{3\times 3\times 3}$, these values are set to $(3,3,3)$, 11, 11, $3C$, and $C$, respectively. OccLinker is integrated on top of this baseline to complement the missing inter-frame motion cues, rather than replacing any component of the original design.
All experiments were conducted using four NVIDIA L20 GPUs with a batch size of 1. We employed the AdamW optimizer on both the SurroundOcc [surroundOcc] and Occ3D [Occ3D] benchmarks, setting the learning rate to $2\times 10^{-4}$, weight decay to $1\times 10^{-2}$, and betas to $(0.9,0.999)$. The overall training took approximately 24 epochs, corresponding to about 1.5 GPU-days in total, including around 1.4 GPU-days for pretraining the base VON and 0.1 GPU-day for fine-tuning OccLinker.
IV-B Evaluation Metrics
Occupancy accuracy metric
For a rigorous and comprehensive evaluation across benchmarks, we employ Intersection over Union (IoU) and Mean Intersection over Union (mIoU) metrics, which are standard in 3D semantic occupancy prediction tasks [PASCAL, Microsoft_COCO, Cityscapes_dataset, Mask_R_CNN]. IoU measures the global overlap between all predicted occupied voxels and their ground-truth counterparts, regardless of semantic categories. It reflects the overall spatial accuracy of the 3D occupancy prediction. In contrast, mIoU is the mean IoU computed by averaging IoUs over all semantic classes, which reflects class-balanced semantic accuracy.
To further evaluate the temporal consistency and occupancy accuracy for both moving and static objects, we adopt the object categorization scheme from Cam4DOcc [Cam4docc], dividing objects into two broad classes: General Moving Objects (GMO) and General Static Objects (GSO). Objects that remain stationary relative to the ground are classified as GSO, whereas those that move relative to the ground belong to the GMO category. This classification scheme enables a fine-grained and quantitative analysis of moving and static objects within the scene. Detailed semantic categories for each class are listed in Table I. The mIoU scores are reported separately for three groups: all classes, GMO classes, and GSO classes.
Classification GMO GSO
Semantic category bicycle, bus, car, construction_vehicle, motorcycle, pedestrian, trailer, truck others, noise, traffic_cone, driveable_surface, other_flat, sidewalk, terrain, manmade, vegetation
Table I: Details of semantic classification. Compared to SurroundOcc, Occ3D includes an additional “others” category under GSO. Both benchmarks also contain an “unoccupied” category, which does not belong to either GMO or GSO.
Temporal consistency metric
To quantify the temporal stability gained by integrating OccLinker with baseline models, we introduce a temporal consistency metric designed to detect and measure changes between consecutive frames. This metric reflects the smoothness and stability of predictions, which is crucial for enhancing the user’s visual experience.
Let $\sigma_{i,n}^{(x,y,z)}$ denote the semantic label of the $n$-th voxel at spatial coordinates $(x,y,z)$ in frame $i$. Define the indicator function $\delta(e_{1},e_{2})=\mathbb{I}(e_{1}\neq e_{2})$, which equals 1 if the two inputs differ and 0 otherwise. Voxel changes between frames $i$ and $j$ fall into two categories: “Moving Object Change” (MOC) and “Static Object Change” (SOC), defined as follows:
Type Condition
MOC $\sigma_{i,n}^{(x,y,z)}\in\text{GMO}\ \text{or}\ \sigma_{j,n}^{(x,y,z)}\in\text{GMO}$
SOC $\sigma_{i,n}^{(x,y,z)}\in\text{GSO}\ \text{and}\ \sigma_{j,n}^{(x,y,z)}\in\text{GSO}$
Table II: Definitions of MOC and SOC voxel changes.
Using these definitions, we define disparity metrics $\Delta_{m}$ and $\Delta_{s}$ to quantify temporal inconsistencies between frames $i$ and $j$, as follows:
$\Delta_{m}(i,j)$ $=\dfrac{1}{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}}\sum\limits_{n=1}^{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right),$ (14)
$\Delta_{s}(i,j)$ $=\dfrac{1}{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}}\sum\limits_{n=1}^{N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}}\delta\left(\sigma_{i,n}^{(x,y,z)},\sigma_{j,n}^{(x,y,z)}\right).$
where $N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mc}}$ and $N_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}sc}}$ represent the number of corresponding voxels in the moving and static regions, respectively.
The overall temporal consistency scores for moving and static objects, $S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}$ and $S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}$, are then computed by aggregating disparities over sequential frames:
$S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m/s}}=1-\dfrac{1}{M-1}\sum\limits_{k=1}^{M-1}\Delta_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m/s}}(k,k+1),$ (15)
where $M$ denotes the total number of frames in the scene. Final consistency scores $\overline{S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}}$ and $\overline{S_{\rm{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}}$ are averaged across all scenes. Higher values indicate smoother and more temporally consistent occupancy predictions.
IV-C Comparison Results
Occupancy accuracy on SurroundOcc benchmark
We compare our method with several state-of-the-art models, including single-frame VON methods such as Atlas [Atlas], TPVFormer [TPVFormer], MonoScene [MonoScene], and SurroundOcc [surroundOcc], as well as history-aware VON methods such as BEVFormer [BEVFormer] and BEVDet4D-Occ [bevdet4d]. To ensure a fair comparison, all baselines and OccLinker are trained on the same ground truth data using identical training protocols. As shown in Table III, integrating OccLinker into existing baselines like MonoScene [MonoScene] and SurroundOcc [surroundOcc] leads to notable performance gains. In particular, the combination of OccLinker with SurroundOcc achieves the best results among all models, improving IoU and mIoU (All) by 1.63% and 0.37%, respectively. Moreover, compared with BEVFormer, which incorporates historical BEV features, our method further boosts performance by 0.41% in IoU and 0.34% in mIoU.
Occupancy accuracy on Occ3D benchmark
We also evaluate our method on the Occ3D benchmark [Occ3D], as shown in Table IV. The baselines include single-frame VON methods such as MonoScene [MonoScene], OccFormer [OccFormer], SurroundOcc [surroundOcc], and ViewFormer [viewformer], as well as history-aware VON methods including SparseOcc [SparseOcc_Liu], BEVDet4D-Occ [bevdet4d], and OPUS-L [opus]. When integrated into single-frame VONs, OccLinker consistently improves the performance of the base models. Furthermore, in ViewFormer+OccLinker, ViewFormer maintains a memory queue in which each frame’s intermediate BEV feature interacts with historical multi-frame BEV representations, while our method further supplements the missing sparse motion features between frames. As a result, ViewFormer+OccLinker surpasses ViewFormer by 0.24% in IoU and 0.94% in mIoU (All).
We note that in Table IV, ViewFormer achieves higher IoU than SurroundOcc, while this advantage is not reflected in $\overline{S_{s}}$. This is mainly because ViewFormer attains higher IoU by aggregating multi-frame BEV features, which captures richer geometric context for individual frames. However, its complex temporal fusion may occasionally propagate redundant or noisy historical information, resulting in less stable transitions between consecutive frames. In contrast, SurroundOcc performs frame-by-frame prediction with limited reliance on long-term history, which naturally leads to smoother temporal behavior and thus higher temporal consistency scores, albeit at the cost of lower per-frame spatial accuracy. Notably, our method improves the temporal consistency of both categories of models by explicitly enhancing sparse inter-frame correlations without introducing additional temporal noise.
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
All GMO GSO
Single-frame VON methods
Atlas [Atlas] 28.66 15.00 12.64 17.35
TPVFormer [TPVFormer] 30.86 17.10 14.04 20.15
MonoScene [MonoScene] 10.04 1.15 0.24 2.07 46.53 81.77
+OccLinker $\Delta$ 13.10 +3.06 1.69 +0.54 0.34 +0.10 3.04 +0.97 54.21 +7.68 83.84 +2.07
SurroundOcc [surroundOcc] 31.49 20.30 18.39 22.20 58.33 91.71
+OccLinker $\Delta$ 33.12 +1.63 20.67 +0.37 18.26 -0.13 23.08 +0.88 60.64 +2.31 92.54 +0.83
History-aware VON methods
BEVDet4D-Occ [bevdet4d] 24.26 14.22 11.10 17.34
BEVFormer [BEVFormer] 30.50 16.75 14.17 19.33 52.84 84.21
+OccLinker $\Delta$ 30.91 +0.41 17.09 +0.34 14.16 -0.01 20.01 +0.68 55.21 +2.37 85.99 +1.78
Table III: Occupancy prediction performance on the SurroundOcc benchmark [surroundOcc]. Performance gains ($\Delta$) are highlighted in bold, and the best results are marked with .
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
All GMO GSO
Single-frame VON methods
MonoScene [MonoScene] 6.06 5.36 6.68
OccFormer [OccFormer] 21.93 21.78 22.06
SurroundOcc [surroundOcc] 51.89 7.24 0.36 13.35 65.35 89.54
+OccLinker $\Delta$ 52.13 +0.24 10.33 +3.09 1.98 +1.62 17.76 +4.41 69.60 +4.25 90.91 +1.37
History-aware VON methods
SparseOcc [SparseOcc_Liu] 30.10
BEVDet4D-Occ [bevdet4d] 39.30 29.09 42.16
OPUS-L [opus] 36.20 31.25 40.44
ViewFormer [viewformer] 70.39 40.46 33.73 46.45 67.26 86.06
+OccLinker $\Delta$ 70.63 +0.24 41.30 +0.84 34.33 +0.60 47.50 +1.05 70.13 +2.87 87.10 +1.04
Table IV: Occupancy prediction performance on the Occ3D benchmark [surroundOcc]. Performance gains ($\Delta$) are highlighted in bold, and the best results are marked with .
Fig. 5: Comparison in a T-junction scenario with partial and dynamic pedestrian occlusions. SurroundOcc+OccLinker produces robust predictions, with the pedestrian being consistently tracked, while SOTA methods show a flickering phenomenon.
Temporal consistency
The $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$ results presented in Tables III and IV show that integrating OccLinker consistently improves the temporal consistency of occupancy predictions across all frames and scenes for all models, highlighting the effectiveness of our approach. This improvement can be attributed to the use of previous static features and the incorporation of motion information, which together provide valuable correlations between static/dynamic cues and occupancy refinement.
IV-D Ablation study
Different combinations of OccLinker
Table V presents the ablation results of different decomposed correlation designs in OccLinker, as illustrated in Fig. 4, when integrated with SurroundOcc [surroundOcc] under $L=1$. M0 denotes the baseline model without OccLinker. M1 removes the correlation between historical and current static features. M2 omits the current static features entirely. M3 excludes the correlation between motion features and current static features. M4 reverses the roles of queries and key/value pairs in the correlation design of Fig. 4. M4 represents the complete version of OccLinker with all components included.
The results in Table V clearly demonstrate that removing any individual stream from the OccLinker module leads to performance degradation in both prediction accuracy and temporal consistency, confirming the necessity of the decomposed correlation modeling. Moreover, the comparison between M4 and M4 highlights that treating previous keyframes and motion features as queries – rather than as keys/values—effectively captures temporally aligned cues essential for robust occupancy refinement.
[…]
M0 31.49 20.30 58.33 91.71
M1 33.04 20.04 60.59 92.25
M2 33.05 19.98 60.09 92.44
M3 32.88 20.10 60.24 92.24
M4 31.97 20.11 60.19 92.01
M4 33.12 +1.63 20.67 +0.37 60.64 +2.31 92.54 +0.83
Table V: Ablation study of decomposed latent correlation modeling in OccLinker when integrated with SurroundOcc. Compared to M4, M4 reverses the roles of queries and key/value pairs. Performance gains over M0 are bolded, and the best results are marked with .
Impact of different types of motion information
Table VI investigates the impact of different types of motion information on SurroundOcc+OccLinker when $L=1$. Specifically, I0 denotes the base model [surroundOcc] without any motion input. I1 uses raw high-frequency intermediate frames as motion input for OccLinker. I2 incorporates optical flow [OpticalFlow], while I3 employs frame difference [Frame_difference] to extract motion cues.
Idx. Motion type IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$ Time $\downarrow$
I0 - 31.49 20.30 58.33 91.71 N/A
I1 Raw Image 32.39 19.45 59.01 91.15 N/A
I2 Optical Flow 32.80 20.27 60.53 92.13 9.221 ms
I3 Frame Diff. 33.12 20.67 60.64 92.54 6.181 ms
Table VI: Ablation study of different motion information types used in OccLinker. The best results are marked with .
Fig. 6: Comparison in a ramp entry scenario with partial occlusions from roadside vegetation. SurroundOcc+OccLinker produces robust and continuous predictions of the drivable surface, while SOTA methods occasionally misclassify the ramp as blocked.
As shown in Table VI, I1 improves IoU over I0, but performs worst in mIoU, $\overline{S_{\rm m}}$, and $\overline{S_{\rm s}}$ among the motion-aware variants. This performance drop is largely due to the presence of redundant and irrelevant information in raw intermediate frames, which complicates the extraction of motion features by OccLinker. Moreover, I3 significantly outperforms I2 in both IoU and mIoU, and achieves slightly better results in $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$. This suggests that frame difference is more effective at capturing abrupt scene changes, such as pedestrians suddenly appearing or vehicles exiting intersections, whereas optical flow may introduce delays in responding to such dynamics. In addition, thanks to the lightweight design of OccLinker, using frame difference further reduces computational overhead by operating on simple differential signals, thereby enhancing inference efficiency.
Impact of different numbers of previous keyframes
We conduct ablation experiments on the temporal context window length $L\in\{0,1,2,3,4\}$ in $\mathcal{W}$. The case of $L=0$ corresponds to SurroundOcc [surroundOcc], which does not utilize any historical keyframes. As shown in Table VII, the best IoU and mIoU are achieved when $L=1$, while the highest $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$ scores occur at $L=2$.
Idx. Window length $L$ IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
L0 0 31.49 20.30 58.33 91.71
L1 1 33.12 20.69 60.64 92.54
L2 2 31.89 20.32 61.00 92.71
L3 3 31.67 20.01 59.01 91.77
L4 4 CUDA Out of Memory
Table VII: Effect of different number of previous keyframes. The best results are marked with .
When introducing historical keyframes (from $L=0$ to $L=1$), both occupancy accuracy metrics (IoU and mIoU) and temporal consistency metrics ($\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$) are improved, indicating that OccLinker enables SurroundOcc to leverage additional historical information and thus enrich temporal context. However, when $L$ is further increased to 2, although temporal consistency continues to improve, larger values of $L$ also introduce more irrelevant or noisy features. This makes it more challenging to associate historical information with the current occupancy prediction and may potentially lead to semantic misalignment.
Moreover, since OccLinker is designed for efficiency, increasing $L$ linearly expands the input volume and computational cost, leading to greater resource consumption during training and inference. Based on these findings, we recommend setting $L$ to 1 or 2 as a balanced choice between performance and efficiency.
Impact of different aggregation strategies in OccLinker
As illustrated in Fig. 4, OccLinker adopts a late aggregation strategy (S2 in Table VIII), where the learned tokens after MHAM are first passed through separate decoders and then aggregated in the occupancy space. In contrast, strategy S1 employs early aggregation by combining the learned tokens in token space before decoding.
The results in Table VIII show that the 3D-space fusion used in S2 yields superior performance. This is because the output of OccLinker functions more effectively as an occupancy correction term when added directly in the occupancy space. Furthermore, the additive nature of S2’s design aligns more naturally with the plug-and-play paradigm than the early fusion approach used in S1.
[…]
S1 Early aggregation 40.58 4,780 MB 0.1598 s
S2 Late aggregation 41.30 4,687 MB 0.1210 s
Table VIII: Effect of different aggregation strategies in ViewFormer+OccLinker. The best results are marked with .
IV-E Case analysis
To qualitatively assess the effectiveness of our method, we compare SurroundOcc+OccLinker with the a single-frame 3D VON baseline (SurroundOcc [surroundOcc]) and a history-aware VON method (BEVDet4D-Occ [bevdet4d]).
Temporal visualization cases
As shown in Fig. 5 (Scene 277, Frames #7–#11), a pedestrian walking along the sidewalk – parallel to the ego vehicle’s trajectory – is intermittently occluded by roadside vegetation. SurroundOcc [surroundOcc] exhibits significant instability, with the pedestrian missing in Frames #7 and #9, highlighting its limitations in temporal modeling. BEVDet4D-Occ [bevdet4d] partially alleviates this issue by fusing historical latent features into the current latent space, but still shows occasional inconsistencies, such as a detection dropout in Frame #8. In contrast, SurroundOcc+OccLinker completely eliminates these flickering artifacts and achieves consistent detection across all occlusion states.
Fig. 7: Several challenging scenarios are presented: pedestrians are partially occluded by vehicles in the first and second columns, and road boundaries appear visually ambiguous in the third column. SurroundOcc+OccLinker yields more accurate predictions, while SOTA methods display noticeable artifacts.
As shown in Fig. 6 (Scene 926, Frames #13-#16), in the ramp entry scenario, the front-right ramp lane in the camera front view is partially occluded by roadside vegetation. SurroundOcc+OccLinker produces robust and continuous predictions of the drivable surface at the junction, while the baseline methods exhibit flickering predictions, occasionally misclassifying the entrance as blocked.
Single-frame visualization cases
Fig. 7 showcases several challenging scenarios: (i) Vehicle-pedestrian occlusion (Scene-0911 Frame #15, Scene-0928 Frame #14): Both SurroundOcc [surroundOcc] and BEVDet4D-Occ [bevdet4d] fail to recover the occupancy state of the partially occluded pedestrian, whereas our method accurately localizes the target with precise geometry. (ii) Ambiguous road boundary (Scene-0923 Frame #28): Baseline methods produce fragmented or erroneous occupancy predictions. In contrast, our approach correctly anticipates the right-turn road geometry and preserves superior shape consistency with the actual road layout.
IV-F Efficiency–performance tradeoff analysis
To ensure a fair comparison, all overhead analysis experiments are conducted on a single NVIDIA L20 GPU. Table Table IX summarizes the results of several state-of-the-art methods, including history-aware approaches (FB-Occ [fb_occ], SparseOcc [SparseOcc_Liu], OPUS-L/M/S/T [opus], and BEVDet4D-Occ [bevdet4d]), the single-frame 3D VON baseline (ViewFormer [viewformer]), and our method combined with ViewFormer. All models are evaluated on the Occ3D benchmark [Occ3D] using the same 2D visual encoder backbone, ResNet-50, which outputs feature maps with a resolution of $(32,88)$. The table reports the mIoU, GPU memory usage during inference, and per-sample latency of each method.
Model mIoU $\uparrow$ Memory (Test) $\downarrow$ Latency $\downarrow$
FB-Occ [fb_occ] 39.11 5,933 MB 0.09 s
SparseOcc [SparseOcc_Liu] 30.10 7,147 MB 0.05 s
OPUS-L [opus] 36.20 10,579 MB 0.16 s
OPUS-M [opus] 35.60 8,912 MB 0.11 s
OPUS-S [opus] 34.20 7,320 MB 0.07 s
OPUS-T [opus] 33.20 6,711 MB 0.03 s
BEVDet4D-Occ [bevdet4d] 39.30 4,689 MB 0.26 s
ViewFormer [viewformer] 40.46 4,582 MB 0.11 s
+OccLinker $\Delta$ 41.30 +0.84 4,687 MB +105 MB 0.12 s +0.01 s
Table IX: Comparison of computational overhead. All models share the same visual encoder ($\Phi_{\rm SE}$) of ResNet-50 backbones. Our result (ViewFormer+OccLinker) in this table is measured for $L=1$. OOM indicates CUDA out of memory. Performance changes ($\Delta$) are highlighted in bold, and best results are marked with .
The results show that ViewFormer+OccLinker achieves state-of-the-art mIoU while maintaining minimal GPU memory usage. The OccLinker plug-in consumes only 105MB, accounting for just 2.29% of ViewFormer’s memory footprint, and introduces a mere 0.01 seconds increase in inference latency. These findings confirm that OccLinker delivers competitive performance with negligible computational overhead.
V Conclusion
This paper proposes OccLinker, a lightweight, plug-and-play module that seamlessly integrates with 3D VONs. OccLinker effectively extracts fine-grained discriminative features by attentively associating historical and motion features with the current feature space through decomposed cross-attention mechanisms. We introduce a new temporal consistency metric to better analyze the effectiveness of OccLinker when combined with SOTA methods in long-term predictions. Extensive visualizations validate the noticeable deflickering effect in several challenging scenarios. Future work will explore the integration of explicit instance-level supervision in occupancy tracking.
Acknowledgments
The study was supported in part by Shenzhen Science and Technology Program under grant number JCYJ20241202130025030.
[Uncaptioned image] Fengcheng Yu received a bachelor’s degree in engineering from the School of Intelligent Engineering of Sun Yat-sen University, majoring in intelligent science and technology, 2025. He is currently pursuing the MS in Computer Science degree with the Viterbi School of Engineering, University of Southern California, California, USA. His research interests includes autonomous driving, high performance computing.
[Uncaptioned image] Haoran Xu received the master’s degree in Software Engineering from Guangdong University of Technology, Guangzhou, China, in 2022. He is currently pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China, and also with Peng Cheng Laboratory, Shenzhen, China. His research interests include multi-agent systems, reinforcement learning, and their applications in connected and autonomous vehicles.
[Uncaptioned image] Canming Xia received the master’s degree in Control Science and Engineering from Guangdong University of Technology, Guangzhou, China, in 2023. Presently, He is pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China. His research interests include decision for autonomous vehicles, reinforcement learning, and representation learning.
[Uncaptioned image] Ziyang Zong received the master’s degree in Computer Science and Technology from Nanjing University of Information Science and Technology, Nanjing, China, in 2023. Presently, she is pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China. Her research interests include domain adaptation, 3D-generation, and representation learning.
[Uncaptioned image] Guang Tan (Member, IEEE) received the Ph.D. degree in computer science from the University of Warwick, Coventry, U.K., in 2007 He was a Professor with the Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences, Shenzhen, China. He has been a Professor with the School of Intelligent Systems Engineering, Sun Yat-sen University, Guangzhou, China, since 2018, where he works on networked systems and machine learning.